OpenStudy (anonymous):

Find the equation of the line passing through (−3, 1) and perpendicular to 3y = −x + 7

OpenStudy (anonymous):

slope of given line line is m=-1/3 if the required line is perpendicular then slope will be m=3 use point slope form y-y0=m(x- x0) where m=3 (x0,y0)=(-3,1)

OpenStudy (anonymous):

would it be y = 3x+10?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

welcome :)